The wavelength of green light from a traffic signal is centered at 5.20 x 10 -5 cm. Calculate the

frequency.


1
Expert's answer
2021-10-20T05:22:47-0400

f = "\frac{c}{λ}"

f = 2.9979 × 10⁸ m/s / 5.20 × 10⁻⁷ m

f = 5.77 × 10¹⁴ Hz